OpenStudy (anonymous):

a student is saving money to buy a used car. the student puts 150 monthly into a savings account that pays 3% annual interest compounded monthly. how long will it take the student to save 5000?

OpenStudy (anonymous):

this is pre calculus 12

OpenStudy (anonymous):

\[FV=R \left[ \left( 1+i \right)^{?}-1 \right]\div i\]

OpenStudy (anonymous):

that is one equation

OpenStudy (anonymous):

another is like \[PV=R \left[ 1-\left( 1+I \right)NEG ^{?}\right]\div i\]

OpenStudy (anonymous):

i meant to put a negative exponent but couldnt do it

OpenStudy (anonymous):

\[A_{1}=150(1+\frac{0.03}{12})\] \[A_{2}=[150(1.0025)](1.0025)\] \[A_{3}=[150(1.0025)^2](1.0025)\] It's a geometric sequence. Use Sum of a G.P \[S_{n}=5000\] \[5000=\frac{ a(r^n-1) }{ r-1 }\]

OpenStudy (anonymous):

its a log and exponent question

OpenStudy (anonymous):

logarithms

OpenStudy (anonymous):

Just work out until youg et to the log part please

OpenStudy (anonymous):

The log part is the easiest part you can get in any log question...

OpenStudy (anonymous):

sub in the values of a and r please. and you have to find n

OpenStudy (anonymous):

can you use the formulas please? i am so confused

OpenStudy (anonymous):

\[5000=\frac{ 150(1.0025)[1.0025^n-1] }{ 0.0025 }\] If you're so lazy to do that, then I will do it for you... \[5000(0.0025)=150(1.0025)[1.0025^n-1]\] \[\frac{5000(0.0025)}{150(1.0025)}=1.0025^n-1\] \[\frac{5000(0.0025)}{150(1.0025)}+1=1.0025^n\]

OpenStudy (anonymous):

Work out that LHS and tell me what you get as a fraction.

OpenStudy (anonymous):

It will take the person 32 months...

OpenStudy (anonymous):

Your equations are dodgy as h*ll.

OpenStudy (anonymous):

It's better to use the formula for Sum of Geometric Progression than what you put up there.

OpenStudy (anonymous):

thanks for your help. believe it or not. i got 32 but was confused because the answer said 2.3 years.. and then i realised it was 32 months converted to years.. derp

OpenStudy (anonymous):

i can do math. i was not being lazy

OpenStudy (anonymous):

just letting you know

OpenStudy (anonymous):

Wouldn't know that you could do math after you put those weird equations up.

OpenStudy (anonymous):

But if you know you can do math, then good job.

OpenStudy (anonymous):

And if the answers say 2.3 years, then it's not 32 months. 32 months equals 2.7 years.....